The volume of water at a given time in a 2000L tank is respesented by the formula V=2000(1- t/45)^2, where t is the time in minutes. Determine the average rate of change in the volume of water in the tank from minute 0 to minute 10, and use it to describe how the volume of water in the tank is changing during that time.
Answer by stanbon(75887)

The volume of water at a given time in a 2000L tank is respesented by the formula V=2000(1- t/45)^2, where t is the time in minutes. Determine the average rate of change in the volume of water in the tank from minute 0 to minute 10, and use it to describe how the volume of water in the tank is changing during that time.
----------------------------------
Average = [V(10)-V(0)]/[10-0]
V(10) = 2000(1-(10/45))^2
V(10) = 2000(1- (2/9)^2 = 2000(7/9)^2 = 1209.876
V(0) = 2000(1-0)^2 = 2000
Average = [1209.88-2000]/[10]
Average = [-790.1234]/10 = -79.01 liters per minute
================================
Cheers,
Stan H.